the wavelength of incident photons is 0.3.. then the wavelength of photon scattered through 60degrees is?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
The ccase will be always in this form:
l-l=lc(1-cosα)
=>l=l+lc(1-cosα)
=>l’=0.3+2.42(1-1/2)
=>l’=1.51pm
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)